OpenStudy (rayford):

x^2-14x+49. what are the 2 binomials

OpenStudy (anonymous):

(x-7)^2

OpenStudy (anonymous):

PLEASE HELP ME! i'M SO LOST AND DON'T KNOW WHAT TO DO! 1. Let f(x) = x2 – 81. Find f–1(x). 2. Let f(x) = x + 2 and g(x) = x2 – 6x + 3. Find g(f(x)). 3. Let f(x) = x – 3 and g(x) = x + 11. Find f(x) • g(x). 4. Let f(x) = x2 – 3x – 7. Find f(–3). 5. Let f(x) = x2 + 2x – 5 and g(x) = x + 6. Find f(x) + g(x). 6. Let f(x) = x2 – 2x + 2 and g(x) = x – 3. Find f(x) • g(x). 7. Let f(x) = –4x + 7 and g(x) = 10x – 6, find f(g(x)). 8. Let f(x) = 5x + 12. Find f–1(x). 9. Let f(x) = x – 2 and g(x) = x2 – 7x – 9. Find f(g(–1)). 10. Let f(x) = 3x2 + x – 3 and g(x) = x2 – 5x + 1. Find f(x) – g(x).